OpenStudy (anonymous):

find the solution set 7x^2+6x-1=0

OpenStudy (denonakavro):

Do you know the quadratic formula, it is the best way

OpenStudy (anonymous):

can you show me how?

OpenStudy (denonakavro):

Check the formula here http://www.purplemath.com/modules/quadform.htm

OpenStudy (zach298):

x=1/7,−1

OpenStudy (er.mohd.amir):

ax^2+bx+c=0 then x={-b+sqrt(b*b-4a*c)}/2a and x={-b-sqrt(b*b-4a*c)}/2a

OpenStudy (denonakavro):

then plug in the a value, b value, and c value and you should get the answers -1, and 1/7

OpenStudy (er.mohd.amir):

7x*x+6x-1=0 7x*x+(7-1)x-1=0 7x*x+7x-x-1=0 7x(x+1)-1(x+1)=0 (x+1)(7x-1)=0

OpenStudy (anonymous):

thanks for the help yall

OpenStudy (denonakavro):

Who medaled zach, it is not even allowed to give direct answers by code of conduct